# Frank–Wolfe algorithm for DC optimization problem

### v2

@@ 3 Piecewise-star-convex functions
The requirement in (1) is formulated with respect to a fixed minimizer $x^{*}$ and must hold for all $x\in{\cal C}$. A milder, yet still powerful, relaxation is to allow the reference minimizer to depend on $x$. In many DC models, the reference minimizer that makes (1) valid for a given $x$ may vary with $x$, which is precisely what piecewise-star-convexity captures. Even this global condition can be too restrictive when $f$ exhibits multiple “basins” of favorable geometry. A natural remedy is to partition the domain into regions within which a local star–type inequality holds with respect to a (local) minimizer:
@@
Let ${\cal C}\subset\mathbb{R}^{n}$ be convex and let ${\cal P}=\{\,{\cal C}_{i}\,:\,i\in I\}$ be a partition of $\cal C$} into nonempty convex subsets with pairwise disjoint interiors and $\bigcup_{i\in I}{\cal C}_{i}={\cal C}$. We say that $f:\mathbb{R}^{n}\to\mathbb{R}$ is piecewise-star-convex on $\cal C$} with respect to $\cal P$} if, for every $i\in I$,
@@
When the partition is trivial (${\cal P}=\{{\cal C}\}$) and $X_{1}^{*}=X^{*}$, Definition 3.3 reduces to Definition 3.1. In practice, piecewise-star-convexity often arises when $f$ is convex on each region ${\cal C}_{i}$ but not across different regions. In this setting, the minimizers $X_{i}^{*}=\arg\min_{{\cal C}_{i}}f$ are global minimizers within their corresponding cells ${\cal C}_{i}$, although they may be only local (and possibly non–global) minimizers of $f$ on the whole domain $\cal C$}. This structure ensures that star–type inequalities hold on each cell (with respect to its minimizers $X_{i}^{*}$), even if such inequalities fail when crossing cell boundaries.
We emphasize that in Definition 3.3, the collection of subsets in item (ii) need not be finite: the piecewise-star-convexity property may still hold when $\cal C$} is covered by infinitely many convex regions. The following example illustrates that finiteness cannot, in general, be assumed.
@@
By construction, $x\in V_{x_{x}^{*}}$ and the whole segment $x_{x}^{*},x]$ lies in $V_{x_{x}^{*}}$. Moreover, $f(y)=|y-x_{x}^{*}|^{2}$ for all $y\in[x_{x}^{*},x]$. Hence, for any $\lambda\in[0,1]$,
@@
We next present a prototypical example of a function that is piecewise-star-convex. This model arises naturally in sparse estimation, robust quadratics, and regularized control. The $\ell_{1}$ penalty acts as a sign–dependent linear perturbation of a positive–definite quadratic. By partitioning $\mathbb{R}^{n}$ into orthants with fixed coordinate signs, the function reduces on each region to a strictly convex quadratic with a unique minimizer. These cell minimizers provide the reference points in Definition 3.3, thereby establishing piecewise-star-convexity even though, globally, such minimizers may be only local.
For use in the examples below and to simplify notation, define
@@
$\Omega_{s}=\bigl{\{}\,x\in\mathbb{R}^{n}:~s_{i}x_{i}\geq 0,\,\text{for }i=1,\ldots,n\bigr{\}},\qquad\qquad\mathbb{R}^{n}=\bigcup_{s\in\{-1,1\}^{n}}\Omega_{s}.$
On each $\Omega_{s}$, the signs of the coordinates are fixed and $\|x\|_{1}=\sum_{i}|x_{i}|=\sum_{i}s_{i}x_{i}=s^{\top}x$, for $x\in\Omega_{s}$. Hence, the restriction of $f$ to $\Omega_{s}$ is the strictly convex quadratic
$f_{s}(x)=\tfrac{1}{2}\,x^{\top}Qx-\beta\,s^{\top}x,\qquad x\in\Omega_{s},$
with constant Hessian $Q$. Since $Q$ is positive definite, $f_{s}$ is strongly convex on $\mathbb{R}^{n}$ with modulus $\lambda_{\min}$, the minimizer $x_{s}^{*}$ is unique. Considering that $f$ coincides with the strictly convex quadratic $f_{s}$ on each orthant $\Omega_{s}$ and $x_{s}^{*}$ is the (unique) minimizer of $f$ on $\Omega_{s}$, the convexity inequality
@@
holds. Therefore $f$ is piecewise–star–convex on $\mathbb{R}^{n}$ with respect to the full partition $\{\Omega_{s}\}_{s\in\{-1,1\}^{n}}$.
Next we preste a constrained DC optimization problem motivated by contrastive learning formulations.
Example 3.6.
Given two disjoint sets of reference points, a set of positive examples $\mathcal{P}=\{\bar{p}_{1},\dots,\bar{p}_{m}\}\subset\mathbb{R}^{n}$ and a set of negative examples $\mathcal{N}=\{\bar{q}_{1},\dots,\bar{q}_{r}\}\subset\mathbb{R}^{n}$, the objective is to identify a point $x$ that lies close to the positive references while remaining far from the negative ones. Here all distances are measured with respect to the standard Euclidean metric. To remain consistent with our general framework, we impose that the search is carried out over a closed and convex feasible set $\mathcal{C}\subset\mathbb{R}^{n}$. Formally, we study the problem
$\min_{x\in\mathcal{C}}f(x),$ (5)
where the Euclidean contrastive objective is defined by
$f(x):=\sum_{i=1}^{m}\lambda_{i}^{+}\,\|x-\bar{p}_{i}\|_{2}^{2}-\sum_{j=1}^{r}\lambda_{j}^{-}\,\|x-\bar{q}_{j}\|_{1},\qquad x\in\mathbb{R}^{n},$ (6)
with $\lambda_{i}^{+}>0$ and $\lambda_{j}^{-}>0$ fixed weights. The squared $\ell_{2}$ terms pull $x$ toward the positive examples, while the negative $\ell_{1}$ term acts as a contrastive regularizer, promoting separation from the negative examples in a coordinate-dependent way. The objective (6) admits a natural DC decomposition of the form
$g(x):=\sum_{i=1}^{m}\lambda_{i}^{+}\,\|x-\bar{p}_{i}\|_{2}^{2},\qquad h(x):=\sum_{j=1}^{r}\lambda_{j}^{-}\,\|x-\bar{q}_{j}\|_{1},$ (7)
where $g$ is smooth and strongly convex, while $h$ is convex but nondifferentiable. Hence similarly to Example 3.5 we can shoe that f=ghf=g-h belongs to the class of piecewise-star-convex DC models. In fact, let $m\geq 1$ and $\lambda_{i}^{+}>0$ for $i=1,\ldots,m$, and define $\Lambda^{+}:=\sum_{i=1}^{m}\lambda_{i}^{+}>0$. Consider $f:\mathbb{R}^{n}\to\mathbb{R}$ given in (6). Then $f$ is piecewise–star–convex on $\mathbb{R}^{n}$ with respect to the partition ${\cal P}=\{{\cal C}_{S}\}_{S}$ induced by the hyperplane arrangement $\{x_{k}=\bar{q}_{j,k}:~j=1,\ldots,r,\ k=1,\ldots,n\}$, where for $S=(s_{1},\ldots,s_{r})$ with $s_{j}\in\{-1,1\}^{n}$,
${\cal C}_{S}:=\bigl{\{}\,x\in\mathbb{R}^{n}:\ s_{j,k}\,(x_{k}-\bar{q}_{j,k})\geq 0,\ \forall j=1,\ldots,r,\ \forall k=1,\ldots,n\,\bigr{\}}.$
Fix $S=(s_{1},\ldots,s_{r})$ with $s_{j}\in\{-1,1\}^{n}$ and consider $x\in{\cal C}_{S}$. For every $j$ and $k$ we have $|x_{k}-\bar{q}_{j,k}|=s_{j,k}(x_{k}-\bar{q}_{j,k})$ on ${\cal C}_{S}$, hence
$\sum_{j=1}^{r}\lambda_{j}^{-}\,\|x-\bar{q}_{j}\|_{1}\;=\;\sum_{j=1}^{r}\lambda_{j}^{-}\sum_{k=1}^{n}s_{j,k}(x_{k}-\bar{q}_{j,k})\;=\;v_{S}^{\top}x-c_{S},$
where $v_{S}:=\sum_{j=1}^{r}\lambda_{j}^{-}\,s_{j}\in\mathbb{R}^{n}$ and $c_{S}:=\sum_{j=1}^{r}\lambda_{j}^{-}\,s_{j}^{\top}\bar{q}_{j}$ is a constant on ${\cal C}_{S}$. Moreover,
$\sum_{i=1}^{m}\lambda_{i}^{+}\,\|x-\bar{p}_{i}\|_{2}^{2}\;=\;\Lambda^{+}\|x\|_{2}^{2}-2\,b^{\top}x\;+\;d,$
with $b:=\sum_{i=1}^{m}\lambda_{i}^{+}\,\bar{p}_{i}$ and $d:=\sum_{i=1}^{m}\lambda_{i}^{+}\,\|\bar{p}_{i}\|_{2}^{2}$. Therefore, on ${\cal C}_{S}$ the function $f$ reduces to the strongly convex quadratic
$f_{S}(x)\;=\;\Lambda^{+}\|x\|_{2}^{2}-(2b+v_{S})^{\top}x\;+\;(d-c_{S}),$
whose Hessian is $2\Lambda^{+}I$ and is positive definite. Since ${\cal C}_{S}$ is closed and convex and $f_{S}$ is strongly convex, $f_{S}$ admits a unique minimizer $x_{S}^{\ast}\in{\cal C}_{S}$. By convexity of $f_{S}$ on the convex set ${\cal C}_{S}$, for all $x\in{\cal C}_{S}$ and all $\lambda\in[0,1]$,
$f\bigl{(}\lambda x_{S}^{\ast}+(1-\lambda)x\bigr{)}\;\leq\;\lambda f(x_{S}^{\ast})+(1-\lambda)f(x).$
This is precisely the star–type inequality required in Definition 3.3 with anchor $x_{S}^{\ast}$ on the cell ${\cal C}_{S}$. Since the interiors of the cells $\{{\cal C}_{S}\}$ are pairwise disjoint and $\bigcup_{S}{\cal C}_{S}=\mathbb{R}^{n}$, the claim follows.
The structure f=ghf=g-h reflects the central principle of contrastive learning: the quadratic component promotes similarity with positive references, while the $\ell_{1}$ component enforces dissimilarity from negative ones. For comprehensive discussions of contrastive learning in the Euclidean setting, see [24]. In the next proposition, we establish that, under a suitable dominance condition on the positive weights, the function $f$ in (6) is coercive and therefore attains at least one global minimizer of (5).
The next example shows how the minimum of convex quadratics induces a polyhedral partition of the space, a construction closely related to Voronoi diagrams and with applications in clustering, facility location, geometry of distance and constrained optimization (see [Libertietall2014] and [1] for a comprehensive survey on this subject). This construction is also connected to Example 3.4.
@@
Next, observe that $f$ admits a DC representation f=ghf=g-h with a smooth convex component $g$ and a convex (typically nonsmooth) component $h$. Indeed,
@@
Therefore, we have f(x)=g(x)h(x)f(x)=g(x)-h(x) with the components given by $g(x):=\tfrac{1}{2}\,x^{\top}Qx)$ and $h(x):=\max_{1\leq i\leq m}\bigl{(}-b_{i}^{\top}x-c_{i}\bigr{)}$. Clearly $g$ is $C^{\infty}$ and convex, while $h$ is the pointwise maximum of affine functions, hence convex, polyhedral, and Lipschitz.
A particularly relevant special case of (8) is obtained by taking pairwise distinct points $a_{1},\dots,a_{m}\in\mathbb{R}^{n}$ and setting $f(x)=\min_{1\leq i\leq m}\,\|x-a_{i}\|^{2},\qquad x\in\mathbb{R}^{n}.$
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be piecewise–star–convex on $\cal C$} with respect to the partition ${\cal P}=\{{\cal C}_{i}\}_{i\in I}$, and let $g:\mathbb{R}^{n}\to\mathbb{R}$ be convex on $\cal C$}. Then $f+g$ is piecewise–star–convex on $\cal C$} with respect to $\cal P$}.
@@
Since $g$ is convex on ${\cal C}_{i}$,
$g\bigl{(}\lambda x_{x}^{*}+(1-\lambda)x\bigr{)}\leq\lambda g(x_{x}^{*})+(1-\lambda)g(x),\quad\forall\,\lambda\in[0,1].$
Summing the inequalities yields the star–type inequality for $f+g$ on ${\cal C}_{i}$ with the same base point $x_{x}^{*}$. As the argument holds for every $i\in I$ and $x\in{\cal C}_{i}$, the claim follows. ∎
@@
If your definition of piecewise–star–convexity explicitly requires the reference points to be minimizers of the function under consideration on each cell, then the statement above remains valid provided $(f+g)$ attains a (local) minimum on each ${\cal C}_{i}$ (e.g., when each ${\cal C}_{i}$ is compact or $(f+g)$ is coercive there). In that case one may choose the reeference points among $\arg\min_{{\cal C}_{i}}(f+g)$.
The next example combines Example 3.5 with Proposition 3.8 to show that a widely used optimization objective is piecewise–star–convex.
Example 3.9.
Let $A\in\mathbb{R}^{n\times n}$ be a full rank column, equivalently, $Q:=A^{\top}A$ a positive definite matrix, let $b\in\mathbb{R}^{n}$, and let $\beta>0$. Consider the function $f:{\mathbb{R}}^{n}\to\mathbb{R}$ defined by
$f(x)=\tfrac{1}{2}\,\|Ax-b\|_{2}^{\,2}-\beta\|x\|_{1},\qquad x\in\mathbb{R}^{n}.$
Expanding the quadratic term gives the equivalent representation
$f(x)\;=\;\tfrac{1}{2}\,x^{\top}Qx-\beta\|x\|_{1}-c^{\top}x\;+\;\tfrac{1}{2}\|b\|_{2}^{2},\qquad Q:=A^{\top}A,\qquad c:=A^{\top}b.$ (9)
Equivalently, write $f=h+g$ with $h(x):=\tfrac{1}{2}\,x^{\top}Qx-\beta\|x\|_{1}$ and $g(x):=-\,c^{\top}x+\tfrac{1}{2}\|b\|_{2}^{2}$. By Example 3.5, $h$ is piecewise–star–convex on the orthant partition when $Q\succ 0$; since $g$ is convex (affine), Proposition 3.8 implies that $f=h+g$ is also piecewise–star–convex on the same partition.
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be piecewise–star–convex on $\cal C$} with respect to the partition ${\cal P}=\{\mathcal{C}_{i}\}_{i\in I}$, where each $\mathcal{C}_{i}$ is closed and convex, interiors are pairwise disjoint, and $\bigcup_{i\in I}\mathcal{C}_{i}=\mathcal{C}$. Let $A\in\mathbb{R}^{n\times n}$ be invertible and $b\in\mathbb{R}^{n}$. Define the affine map
@@
the transformed domain and cells $\tilde{\mathcal{C}}:=A^{-1}(\mathcal{C}-b)$ and $\tilde{\mathcal{C}}_{i}:=A^{-1}(\mathcal{C}_{i}-b)$ and the function $\tilde{f}:\tilde{\mathcal{C}}\to\mathbb{R}$ by
$\tilde{f}(z):=f(Az+b),\qquad z\in\tilde{\mathcal{C}}$.
@@
(i) Since $A$ is invertible, $\Psi$ is a bijective affine homeomorphism with inverse $\Psi^{-1}(x)=A^{-1}(x-b)$. If $z\in\tilde{\mathcal{C}}$ then $x:=\Psi(z)=Az+b\in\mathcal{C}$, so $\tilde{f}(z)=f(x)$ is well defined. Because $A^{-1}$ and the translation $x\mapsto x-b$ are continuous linear/affine maps, each $\tilde{\mathcal{C}}_{i}=A^{-1}(\mathcal{C}_{i}-b)$ is closed and convex; interiors remain pairwise disjoint and $\bigcup_{i}\tilde{\mathcal{C}}_{i}=\tilde{\mathcal{C}}$. Thus $\tilde{\mathcal{P}}$ is a partition of $\tilde{\mathcal{C}}$.
(ii) For each $i\in I$, let $X_{i}^{*}:=\arg\min_{\mathcal{C}_{i}}f$ and $\tilde{X}_{i}^{*}:=\arg\min_{\tilde{\mathcal{C}}_{i}}\tilde{f}$. We claim
@@
(iii) Fix $i\in I$ and $z\in\tilde{\mathcal{C}}_{i}$, and set $x:=\Psi(z)=Az+b\in\mathcal{C}_{i}$. Because $f$ is piecewise–star–convex on $\mathcal{C}_{i}$, there exists $x_{x}^{*}\in X_{i}^{*}$ such that, for all $\lambda\in[0,1]$,
@@
Define $z_{z}^{*}:=A^{-1}(x_{x}^{*}-b)$. By (10), $z_{z}^{*}\in\tilde{X}_{i}^{*}$. Using the affinity of $\Psi$,
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be locally Lipschitz on a convex set $\cal C$}. Assume $f$ is piecewise–star–convex on $\cal C$} with respect to the partition ${\cal P}=\{{\cal C}_{i}\}_{i\in I}$ of Definition 3.3. Fix $i\in I$ and $x\in{\cal C}_{i}$. Then there exists $x_{x}^{*}\in X_{i}^{*}:=\arg\min_{{\cal C}_{i}}f$ (a minimizer on the cell ${\cal C}_{i}$; hence, a local minimizer for $f$ on $\cal C$}) such that
@@
Subtract $f(x)$, divide by $\lambda>0$, and let $\lambda$ goes to 0 to obtain $f^{\circ}(x;\,x_{x}^{*}-x)\ \leq\ f(x_{x}^{*})-f(x).$ By the Clarke calculus (e.g., Theorem 2.1), $f^{\circ}(x;d)=\max_{v\in\partial_{c}f(x)}v^{\top}d$, hence $v^{\top}(x_{x}^{*}-x)\leq f^{\circ}(x;\,x_{x}^{*}-x)$ for all $v\in\partial_{c}f(x)$, which proves (12). ∎
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be locally Lipschitz and piecewise-star-convex on $\cal C$} with respect to the partition ${\cal P}=\{{\cal C}_{i}\}_{i\in I}$. Suppose that $\bar{x}\in{\cal C}$ is a Clarke stationary point of the constrained problem $\min_{x\in{\cal C}}f(x),$ that is,
$v^{\top}(x-\bar{x})\ \geq\ 0\qquad\forall\,v\in\partial_{c}f(\bar{x}),\ \forall\,x\in{\cal C}.$ (13)
If, in addition, $\bar{x}\in{\cal C}_{i_{0}}$ for some $i_{0}\in I$, then $\bar{x}$ is a global minimizer of $f$ on the corresponding cell ${\cal C}_{i_{0}}$, namely, $\bar{x}\in X_{i_{0}}^{*}:=\arg\min_{x\in{\cal C}_{i_{0}}}f(x).$ In particular, this means that $f(\bar{x})=\min_{x\in{\cal C}_{i_{0}}}f(x).$
@@
Apply Proposition 3.11 with $x=\bar{x}$ and $i=i_{0}$ to obtain some $x_{\bar{x}}^{*}\in X_{i_{0}}^{*}$ with
$f(x_{\bar{x}}^{*})-f(\bar{x})\ \geq\ v^{\top}(x_{\bar{x}}^{*}-\bar{x})\qquad\forall\,v\in\partial_{c}f(\bar{x}).$
By stationarity (13) (taking $x=x_{\bar{x}}^{*}\in{\cal C}_{i_{0}}\subset{\cal C}$), the right-hand side is $\geq 0$, hence $f(x_{\bar{x}}^{*})\geq f(\bar{x})$. Since $x_{\bar{x}}^{*}\in X_{i_{0}}^{*}$ minimizes $f$ on ${\cal C}_{i_{0}}$, we also have $f(x_{\bar{x}}^{*})\leq f(\bar{x})$. Therefore $f(x_{\bar{x}}^{*})=f(\bar{x})$ and $\bar{x}\in X_{i_{0}}^{*}$. ∎

### v3

@@ 4 Piecewise star-convexity with a nonsmooth DC structure
The condition in (6) is stated with respect to a fixed minimizer $x^{\ast}$ and must hold for all $x\in{\cal C}$. A milder relaxation is to allow the reference minimizer to depend on $x$, a behavior that often arises in DC models and is captured by piecewise star-convexity. When $f$ has several “basins” of favorable geometry, even this can be too restrictive globally, so we partition the domain into regions where a star-type inequality holds relative to a local minimizer:
@@
Let ${\cal C}\subset\mathbb{R}^{n}$ be convex and let ${\cal P}=\{\,{\cal C}_{i}\,:\,i\in I\}$ be a partition of $\cal C$} into nonempty convex subsets and $\bigcup_{i\in I}{\cal C}_{i}={\cal C}$. We say that $f:\mathbb{R}^{n}\to\mathbb{R}$ is piecewise-star-convex on $\cal C$} with respect to $\cal P$} if, for every $i\in I$,
@@
When the partition is trivial, that is, ${\cal P}=\{{\cal C}\}$ and $X_{1}^{*}=X^{*}$, Definition 4.3 reduces to Definition 4.1. In many applications, piecewise star-convexity arises because $f$ has a favorable geometry on each region ${\cal C}_{i}$ (for instance, it may be convex or star-convex when restricted to ${\cal C}_{i}$) but this geometry does not persist across different regions. Accordingly, the sets $X_{i}^{*}=\arg\min_{{\cal C}_{i}}f$ consist of global minimizers of the cellwise problems, yet such points need not be global minimizers of $f$ on $\cal C$} and may fail to be even local minimizers of $f$ on $\cal C$}. The key point is that the star-type inequalities in Definition 4.3 are enforced within each cell relative to $X_{i}^{*}$, while no compatibility is required when one crosses cell boundaries.
We emphasize that, in Definition 4.3, the collection of subsets in item (ii) need not be finite. Piecewise star-convexity may still hold when $\cal C$} is covered by infinitely many convex regions. The following example exhibits an infinite partition and an objective function with a nonsmooth difference-of-convex structure.
@@
By construction, $x\in V_{x_{x}^{*}}$ and the whole segment between $x_{x}^{*}$ and $x$ lies in $V_{x_{x}^{*}}$. Moreover, $f(y)=|y-x_{x}^{*}|^{2}$ for all $y$ belonging the segment between $x_{x}^{*}$ and $x$. Hence, for any $\lambda\in[0,1]$,
@@
We next present a prototypical example of a piecewise star-convex function with a nonsmooth DC structure. This model arises naturally in sparse estimation, robust quadratic fitting, and regularized control. The $\ell_{1}$ term is convex but nonsmooth, and it induces a sign-dependent linear perturbation of a positive-definite quadratic. By partitioning $\mathbb{R}^{n}$ into orthants with fixed coordinate signs, the objective reduces on each region to a strictly convex quadratic with a unique minimizer. These cell minimizers provide the reference points in Definition 4.3, which establishes piecewise star-convexity even though, globally, the problem remains nonconvex. For use in the examples below and to simplify notation, define
@@
$\Omega_{s}=\bigl\{\,x\in\mathbb{R}^{n}:~s_{i}x_{i}\geq 0\text{ for }i=1,\ldots,n\bigr\},\qquad\qquad\mathbb{R}^{n}=\bigcup_{s\in\{-1,1\}^{n}}\Omega_{s}.$
On each $\Omega_{s}$ the signs of the coordinates are fixed and $\|x\|_{1}=\sum_{i}|x_{i}|=\sum_{i}s_{i}x_{i}=s^{\top}x$ for $x\in\Omega_{s}$. Hence the restriction of $f$ to $\Omega_{s}$ is the strictly convex quadratic
$f_{s}(x)=\tfrac{1}{2}\,x^{\top}Qx-\beta\,s^{\top}x,\qquad x\in\Omega_{s},$ (10)
with constant Hessian $Q$. Since $Q$ is positive definite, $f_{s}$ is strongly convex on $\Omega_{s}$ with modulus $\lambda_{\min}(Q)$; hence the minimizer $x_{s}^{*}$ is unique. Considering that $f$ coincides with the strictly convex quadratic $f_{s}$ on each orthant $\Omega_{s}$ and $x_{s}^{*}$ is the (unique) minimizer of $f$ on $\Omega_{s}$, the convexity inequality
@@
holds. Therefore $f$ is piecewise–star–convex on $\mathbb{R}^{n}$ with respect to the full partition $\{\Omega_{s}\}_{s\in\{-1,1\}^{n}}$. A particularly interesting case is when $Q=\alpha I_{n}$ with $\alpha>0$. In this case, it follows from (10) that on each orthant $\Omega_{s}$ we have $f_{s}(x)=\tfrac{\alpha}{2}\,\|x\|_{2}^{2}-\beta\,s^{\top}x,$ whose unique minimizer on $\Omega_{s}$ solves $\alpha x-\beta s=0$, hence $x_{s}^{*}=(\beta/\alpha)\,s\in\Omega_{s}.$ Therefore all $2^{n}$ points $(\beta/\alpha)\,s$, with $s\in\{-1,1\}^{n}$, are global minimizers of the function $f$, and the minimum value is $\min_{x\in\mathbb{R}^{n}}f(x)=-\,{n\beta^{2}}/{(2\alpha)},$ which is independent of $s$.
The next example shows how the minimum of convex quadratics induces a polyhedral partition of the space, a construction closely related to Voronoi diagrams and with applications in clustering, facility location, geometry of distance and constrained optimization, see [Libertietall2014] and [2] for a comprehensive survey on this subject. This construction is also connected to Example 4.4.
@@
Next, observe that $f$ admits a nonsmooth DC representation f=ghf=g-h with a smooth convex component $g$ and a convex (typically nonsmooth) component $h$. Indeed,
@@
Therefore, we have f(x)=g(x)h(x)f(x)=g(x)-h(x) with the components given by $g(x):=\tfrac{1}{2}\,x^{\top}Qx$ and $h(x):=\max_{1\leq i\leq m}\bigl(-b_{i}^{\top}x-c_{i}\bigr)$. Clearly $g$ is $C^{\infty}$ and convex, while $h$ is the pointwise maximum of affine functions, hence convex, polyhedral, and Lipschitz.
A particularly relevant special case of (11) is obtained by taking pairwise distinct points $a_{1},\dots,a_{m}\in\mathbb{R}^{n}$ and setting $\varphi_{i}(x)=\tfrac{1}{2}(x-a_{i})^{\top}Q(x-a_{i})$. In this case, each cell $\mathcal{C}_{i}$ contains $a_{i}$, and $\min_{x\in\mathcal{C}_{i}}f(x)=\min_{x\in\mathcal{C}_{i}}\varphi_{i}(x)=\varphi_{i}(a_{i})=0,$ with the minimizer in $\mathcal{C}_{i}$} given by $x_{i}^{\ast}=a_{i}$. Therefore, all cells share the same minimum value 0.
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be piecewise-star-convex on $\cal C$} with respect to the partition ${\cal P}=\{{\cal C}_{i}\}_{i\in I}$, and let $g:\mathbb{R}^{n}\to\mathbb{R}$ be convex on $\cal C$}. Then $f+g$ is piecewise-star-convex on $\cal C$} with respect to $\cal P$}.
@@
Since $g$ is convex on ${\cal C}_{i}$, we have $g\bigl(\lambda x_{x}^{*}+(1-\lambda)x\bigr)\leq\lambda g(x_{x}^{*})+(1-\lambda)g(x)$, for all $\lambda\in[0,1]$. Summing these inequalities yields the star–type inequality for $f+g$ on ${\cal C}_{i}$ with the same base point $x_{x}^{*}$. As the argument holds for every $i\in I$ and $x\in{\cal C}_{i}$, the claim follows. ∎
@@
The Example 3.2 follows immediately from Example 4.5 and Proposition 4.7.
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be piecewise–star–convex on $\cal C$} with respect to the partition ${\cal P}=\{\mathcal{C}_{i}\}_{i\in I}$, where each $\mathcal{C}_{i}$ is nonempty convex and $\bigcup_{i\in I}\mathcal{C}_{i}=\mathcal{C}$. Let $A\in\mathbb{R}^{n\times n}$ be invertible and $b\in\mathbb{R}^{n}$. Define the affine map $\Psi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ by
@@
the transformed sets $\tilde{\mathcal{C}}:=A^{-1}(\mathcal{C}-b)$ and $\tilde{\mathcal{C}}_{i}:=A^{-1}(\mathcal{C}_{i}-b)$ and the function $\tilde{f}:\tilde{\mathcal{C}}\to\mathbb{R}$ by
$\tilde{f}(z):=f(\Psi(z)),\qquad z\in\tilde{\mathcal{C}}$.
@@
Since the matrix $A$ is invertible, the function $\Psi$ is a bijective affine homeomorphism with inverse $\Psi^{-1}(x)=A^{-1}(x-b)$. If $z\in\tilde{\mathcal{C}}$ then $x:=\Psi(z)=Az+b\in\mathcal{C}$, so $\tilde{f}(z)=f(x)$ is well defined. Because $A^{-1}$ and the translation $x\mapsto x-b$ are continuous linear/affine maps, each $\tilde{\mathcal{C}}_{i}=A^{-1}(\mathcal{C}_{i}-b)$ is convex and $\bigcup_{i}\tilde{\mathcal{C}}_{i}=\tilde{\mathcal{C}}$. Thus $\tilde{\mathcal{P}}$ is a partition of $\tilde{\mathcal{C}}$. In addition, for each $i\in I$, let $X_{i}^{*}:=\arg\min_{\mathcal{C}_{i}}f$ and $\tilde{X}_{i}^{*}:=\arg\min_{\tilde{\mathcal{C}}_{i}}\tilde{f}$. We claim
@@
Now, fix $i\in I$ and $z\in\tilde{\mathcal{C}}_{i}$, and set $x:=\Psi(z)=Az+b\in\mathcal{C}_{i}$. Because $f$ is piecewise-star-convex on $\mathcal{C}_{i}$, there exists $x_{x}^{*}\in X_{i}^{*}$ such that, for all $\lambda\in[0,1]$,
@@
Define $z_{z}^{*}:=A^{-1}(x_{x}^{*}-b)$. By (12), $z_{z}^{*}\in\tilde{X}_{i}^{*}$. Using the affinity of $\Psi$, we conclude that
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be locally Lipschitz on a convex set $\cal C$}. Assume $f$ is piecewise–star–convex on $\cal C$} with respect to the partition ${\cal P}=\{{\cal C}_{i}\}_{i\in I}$ of Definition 4.3. Fix $i\in I$ and $x\in{\cal C}_{i}$. Then there exists $x_{x}^{*}\in X_{i}^{*}:=\arg\min_{{\cal C}_{i}}f$ a minimizer on the cell ${\cal C}_{i}$ such that
@@
Subtract $f(x)$, divide by $\lambda>0$, and let $\lambda$ goes to 0 to obtain $f^{\circ}(x;\,x_{x}^{*}-x)\ \leq\ f(x_{x}^{*})-f(x).$ By the Clarke calculus, e.g., Theorem 2.1, $f^{\circ}(x;d)=\max_{v\in\partial_{c}f(x)}v^{\top}d$, hence $v^{\top}(x_{x}^{*}-x)\leq f^{\circ}(x;\,x_{x}^{*}-x)$ for all $v\in\partial_{c}f(x)$, which proves (14). ∎
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be locally Lipschitz and piecewise-star-convex on $\cal C$} with respect to the partition ${\cal P}=\{{\cal C}_{i}\}_{i\in I}$. Suppose that $\bar{x}\in{\cal C}$ is a Clarke stationary for problem (1), i.e., there exists $v\in\partial_{c}f(\bar{x})$ such that
$v^{\top}(x-\bar{x})\geq 0\qquad\forall\,x\in{\cal C}.$ (15)
In addition, for any $i_{*}\in I$ such that $\bar{x}\in{\cal C}_{i_{*}}$, then $\bar{x}$ is a global minimizer of $f$ on the corresponding cell ${\cal C}_{i_{*}}$, namely, $\bar{x}\in X_{i_{*}}^{*}:=\arg\min_{x\in{\cal C}_{i_{*}}}f(x).$ In particular, this means that $f(\bar{x})=\min_{x\in{\cal C}_{i_{*}}}f(x).$
@@
Apply Proposition 4.9 with $x=\bar{x}$, $v\in\partial_{c}f(\bar{x})$ and $i=i_{*}$ to obtain some $x_{\bar{x}}^{*}\in X_{i_{*}}^{*}$ with
$f(x_{\bar{x}}^{*})-f(\bar{x})\geq v^{\top}(x_{\bar{x}}^{*}-\bar{x}).$
By condition (15), taking $x=x_{\bar{x}}^{*}\in{\cal C}_{i_{*}}\subset{\cal C}$, the right-hand side is nonnegative, hence we conclude that $f(x_{\bar{x}}^{*})\geq f(\bar{x})$. Since $x_{\bar{x}}^{*}\in X_{i_{*}}^{*}$ minimizes $f$ on ${\cal C}_{i_{*}}$, we also have $f(x_{\bar{x}}^{*})\leq f(\bar{x})$. Therefore, we have $f(x_{\bar{x}}^{*})=f(\bar{x})$, which implies that $\bar{x}\in X_{i_{*}}^{*}$. ∎